OpenStudy (anonymous):

The sum of two numbers, x and y, is 12. The difference of x and two times y is 6. What are the values of x and y?

OpenStudy (anonymous):

The sum of two numbers, x and y, is 12. The difference of x and two times y is 6. What are the values of x and y? ----------------- Equations: x + y = 12 x - y = 6 ------------ Add and solve for "x": 2x + 18 x = 9 ---- y = 12-x = 3

OpenStudy (anonymous):

is that the answer bro

OpenStudy (anonymous):

the choices are A. x=8,y=4 B.x=10,y=4 C.x=18,y=-6 D.x=20,y=-8

OpenStudy (anonymous):

@elic2k15

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

y = 12-x = 3

OpenStudy (anonymous):

B.is actually x=10,y=2 but do you know the answer from the choices above

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

you know the answerr??

OpenStudy (anonymous):

@elic2k15 please give me the answerrr !!!!!!!!!!!

OpenStudy (anonymous):

@posiimage please give me the answer to this equation i beg of you